# Enhancing Ultra High Resolution Remote Sensing Imagery Analysis with ImageRAG

### v2

@@ X Conclusion and Future Work
In the future, we plan to apply more RAG techniques to ImageRAG, particularly enhancing the ranking component. We will also prioritize optimizing ImageRAG’s efficiency and scalability. For the inferring model, we will investigate how to minimize the negative influence of inaccurate visual cues. Moreover, to boost performance on specialized imaging data, we intend to introduce adapter module with trainable parameters. This will enable effective adaptation to domain-specific data such as SAR, infrared, and hyperspectral imagery. The approach may adopt an in-context learning style, minimizing the need for extensive additional training data while preserving the model’s general capabilities.

### v3

@@ XIII Conclusion and Future Work
In the future, we plan to apply more RAG techniques to ImageRAG, particularly enhancing the ranking component. We will also prioritize optimizing ImageRAG’s efficiency and scalability. For the inferring model, we will investigate how to minimize the negative influence of inaccurate visual cues. Moreover, to boost performance on specialized imaging data, we intend to introduce adapter module with trainable parameters. This will enable effective adaptation to domain-specific data such as SAR, infrared, and hyperspectral imagery [131]. The approach may adopt an in-context learning style, minimizing the need for extensive additional training data while preserving the model’s general capabilities.
XIV Acknowledgement
This work is supported by National Key R&D Program of China (2024YFB3908400).